# Provably Robust Multi-bit Watermarking for AI-generated Text

### v3

@@ 5.1 Experiment setup
Metrics and evaluation methods. We use match rate to measure the proportion of generated text that can exactly extract embedded watermark message without error. Several previous work [40, 66, 67, 68] in watermarking adopted another metric: bit accuracy (the ratio of correctly extracted message bits). For the comparison of these 2 metrics, see Appendix F. We use edit distance to quantify the attacker’s change to the watermarked text. Edit distance (Levenshtein distance) [44] measures the difference between two texts defined as the minimum number of single-token edits (insertions, deletions, or substitutions) required to change one text into the other. Additionally, to measure theoretical robustness, we define the metric robust bound for a dataset as the average robust bound for each generated text. The robust bound for each paragraph is calculated using Algorithm 5 in Appendix C. We set the error probability threshold in Definition 4.2 as $\alpha=10^{-3}$. For text quality, following the work in [29], we use perplexity (PPL) computed by a large oracle model GPT-3.5 (text-davinci-003) [50].
5.2 Main results
@@
Table 2: Compare our method with existing baselines, on match rate and extraction time (per text paragraph).
@@
Match rate (%) Time (s) Match rate (%) Time (s) Match rate (%) Time (s) Match rate (%) Time (s) Match rate (%) Time (s)
Fernandez et al. [17] 99.6 0.12 99.6 0.34 99.2 5.04 98.0 110 NA 29000 (Estimated)
Wang et al. [63] 99.6 0.16 98.8 0.58 98.4 3.14 97.2 35.5 NA 8300 (Estimated)
Yoo et al. [68] 86.4 0.01 73.6 0.01 49.2 0.01 30.4 0.01 8.4 0.01
Cohen et al. [12] 93.2 0.01 88.8 0.02 78.4 0.02 65.6 0.03 27.2 0.04
Ours 98.8 0.04 98.0 0.06 97.6 0.10 96.0 0.18 94.0 0.6
@@
Our method maintains the quality of texts generated by LLMs. We also explore the influence of our watermark on the quality of text generated by an LLM. Experiments are conducted on OpenGen dataset using 3 different models—Llama2, Guanaco, Falcon, and different bit lengths of 12, 16, 20, 24, 28, and 32. Other hyperparameters are under the default settings. Figure 6 uses box plots to compare the perplexity distribution of human-written text with that of texts generated by three different models (Llama2, Guanaco, and Falcon), both with and without watermarking. Note that a lower perplexity indicates higher text quality. The results indicate that watermarking only slightly harms the quality of AI-generated text, but the effect is minor. Prior to the application of our watermark, the perplexity values associated with text generated by Falcon predominantly fall within the interval of $[6,10]$. Following the integration of our watermark, a very slight elevation in perplexity occurs, with the generated text now primarily falling within the range of $[6.5,10.5]$.
Figure 11 in Appendix B shows that the change of embedded bit length has negligible impact on text quality for all models, as indicated by the almost flat curves. These results suggest that our watermarking method subtly affects the quality of the generated text, ensuring the utility of watermarked AI-generated content in practical applications.
Our method outperforms baselines [68, 17, 63]. To compare our approach with baselines [68, 17, 63, 12] 222 [12] did not provide source code, we implemented their algorithm by ourselves., we measure match rate and extraction time with bit lengths set to 12, 16, 20, 24, 32 separately, while other parameters are kept at their default values. We do not compare with  [6] since they are not open-sourced. Additionally, their method is not practical when embedding 20-bit messages, see our analysis in Appendix E.
@@
Figure 7: (a) compares adopting Reed-Solomon code and BCH code. (b) compares with and without adopting balanced segment assignment.
Impact of using other ECC schemes. We also explore the impact of using ECC other than Reed-Solomon code in our watermark, specifically switching to the widely adopted binary BCH code [7]. Experiments are conducted on the default model and dataset, for bit lengths of 12, 16, 20, 24, 28, and 32, with other hyperparameters at their default settings. Figure 7a shows the results of this experiment. We observe that using Reed-Solomon consistently results in a higher match rate across all cases. More importantly, as the bit length increases, the performance gap between Reed-Solomon and BCH codes becomes significant. For example, when $b=12$, using Reed-Solomon code achieves match rate of $98.8\%$ while using BCH achieves $95.2\%$. But when $b=32$, the match rate of using Reed-Solomon code drops about five percent to $94.0\%$ while using BCH code drops about $35\%$ percent to $60.0\%$. This demonstrates that Reed-Solomon code is much more suitable for our watermarking scheme than BCH code. The essential reason is that binary BCH codes excel at correcting errors in individual bits, while Reed-Solomon codes are more effective at correcting burst errors. Our scheme packs multiple bits into one segment and embeds each segment into a group of tokens. Thus, when errors occur during decoding, they often take the form of burst errors, where only one or two segments are incorrect, but multiple bits within those segments may be wrong. This makes Reed-Solomon a better choice for our scenario compared to BCH.
Impact of adopting balanced segment assignment. To explore the effect of our designed balanced segment assignment algorithm, we compare it with the original segment assignment algorithm. Experiments are conducted on the default model and dataset, for bit lengths of 20, 24, 28, 32 and 36, with other hyperparameters under the default settings. Figure 7b shows the results of this experiment. We observe that using balanced segment assignment consistently results in a higher match rate across all cases. Furthermore, as the bit length increases, the performance gap between adopting balanced segment assignment and the original segment assignment becomes significant. For example, when $b=20$, the gap is $2\%$. But when $b=32$, the gap raises to about $5\%$. This demonstrates the effectiveness of balanced segment assignment in improving the performance of our scheme. Making the token assignment for different segments more balanced reduces the probability of certain segment being allocated fewer tokens and therefore being decoded incorrectly. This enhances the accuracy of watermark extraction.
5.3 Impact of hyperparameters
@@
Impact of model sizes and LLM decoding algorithms. To further showcase the generalizability of our method, we conduct experiments across different model sizes of Llama2 model (7B, 13B, and 70B) with other hyperparameters under default settings. We also conduct experiments using different LLM decoding algorithms including greedy sampling, multinomial sampling, and nucleus sampling. For the temperature in sampling, we use 0.7 following prevalent works in LLMs [49]. The results are presented in Table 3 and Table 4 in Appendix B. Our method yields consistently similar performance across these varied conditions, indicating that it possesses robust generalization capabilities.
Impact of $T$. To demonstrate the impact of text generation length $T$ on the match rate and robustness, we conduct experiments with $T$ set at 100, 150, 200, 250 and 300 tokens, while all other parameters are kept at their default settings. Figure 9 in Appendix B presents the outcomes of these experiments. As can be observed from Figure 9a, with all other hyperparameters held constant, an increase in $T$ generally leads to a higher match rate. This effect can be attributed to the fact that a larger $T$ value results in a higher average number of tokens allocated to each segment, which reduces the error probability on each segment since the probability that the correct candidate has the most green tokens increases.
Additionally, Figure 9b shows that a larger $T$ enhances robustness. In accordance with the conclusion of Theorem 4.1, under the same $\eta$, the RHS of Inequality 5 decreases with the increase of $T$. Therefore the robust bound will increase with the increase of $T$.
Impact of $\delta$. To demonstrate the impact of bias $\delta$ on match rate, robustness, and text quality (measured by perplexity), we conduct experiments with $\delta$ ranging from 2 to 10, while keeping other hyperparameters at their default settings. The results are presented in Figure 10 in Appendix B. It can be observed that as $\delta$ increases, both match rate and robustness improve, but text quality deteriorates. Theoretically, when other hyperparameters are fixed, an increase in $\delta$ increases the total number of green tokens for each segment. This further decreases the probability of extraction errors on each segment, because for each segment, the probability for incorrect value corresponding to max green tokens decreases. Thus when $\delta$ is relatively small, increasing $\delta$ improves match rate. On the other hand, when $\delta$ increases, the green token number assigned to each segment d will increase by expectation. This will increase $p^{(n)}_{0}$ in Theorem 4.1. Thus, the RHS of Inequality 5 decreases, thereby increasing the robust bound. However, with the increase of $\delta$, the introduced bias in tokens sampling also becomes stronger and the text quality degrades, which is consistent with the observation in [29]. In conclusion, the change in $\delta$ brings a tradeoff between correctness, robustness, and text quality.
@@ 6 Conclusion
Discussion: Research Ethics
Our research focuses on enhancing the auditing of LLM generated content, preventing and tracing malicious LLM users generating content like fake news and plagiarism. We strictly follow ethical guidelines, avoiding deceptive practices, unauthorized disclosures, live system experiments, or any actions that could jeopardize the well-being of our team. Consequently, our work does not introduce new ethical concerns beyond those already associated with LLMs.
Discussion: Open Science Policy
Our study adheres to open science principles and fully supports artifact evaluation by guaranteeing the availability, functionality, and reproducibility of our work. We anonymously share our code at https://github.com/randomizedtree/segment-watermark. We are committed to submitting our work for artifact evaluation.
@@
Appendix B Some Experimental Results
In this section, we show some experiment results not included in the main body of the paper due to space limitations. Figure 9 shows the match rate and robust bound of watermarked text under different $T$. Figure 10 shows the match rate, robust bound, perplexity of watermarked text under different $\delta$. Figure 11 shows the perplexity of texts generated by different LLMs across different bit lengths. Table 3 shows the match rate and robust bound of watermarked text for LLMs with different sizes. Table 4 shows the match rate and robust bound of watermarked text generated by different LLM decoding algorithms. Table 5 shows examples of watermarked text.
Figure 9: Impact of text length $T$ on match rate and robust bound.
Figure 10: Impact of $\delta$ on match rate, robust bound, and perplexity.
Table 3: Match rate and robust bound of our method for LLMs with different sizes.
Model size LLaMA-2-7b LLaMA-2-13b LLaMA-2-70b
Match rate 97.6 98.4 98.0
Robust bound 17.3 16.9 17.2
Table 4: Impact of LLM decoding algorithm on match rate and robust bound.
Greedy sampling Multinomial sampling Nucleus sampling
Match rate 97.6 98.0 97.2
Robust bound 17.3 17.5 16.8
Table 5: Examples of our watermark embedding and extraction results. The content in the "Prompt" column comes from datasets we use. The "Real Completion" column represents the ground truth completion from datasets. The "Watermarked Text" column represents text embedded with watermark information through our watermarking algorithm. The "Orig bits" column stands for the original message embedded in the watermarked text, whereas the "Extr bits" column represents the message extracted from the watermarked text using our watermarking algorithm.
Prompt Real Completion Watermarked Text Orig bits Extr bits
… In the heavy fighting that followed, Kokusho and around 100 of his men were killed, ending that attack. Meanwhile, Kawaguchi’s 2nd Battalion, under Major Masao Tamura, assembled for their planned assault against Hill 80 from the jungle south of the ridge. Marine observers spotted Tamura’s preparations and called in artillery fire. At about 22:00, a barrage from twelve guns hit Tamura’s position. In response, two companies of Tamura’s troops numbering about 320 men charged up Hill 80 with fixed […continues] At 03:40, artillery shells landed among Tamura’s troops, signaling Keating’s artillery concentrations and starting their attack off late. As the Marines concentrated artillery fires on his men, their commanding officer realized they would not make the ridge […continues] 11000 10011 01000 00111 11000 10011 01000 00111
… The ship was scheduled to leave for Australia on Monday. The last time there was a norovirus outbreak on the ship was back in 2012. According to Yahoo, health officials conducted a series of tests, and they confirmed that the ill- -ness was norovirus, but the outbreak does seem to be going away. The norovirus usually lasts for one to three days, and those infected may experience stomach pains, vomiting, diarrhea and nausea. Princess Cruises released a statement saying that those who were infected were […continues] -nesses were caused by the contagious norovirus, a highly contagious viral gastroenteritis that can cause diarrhea, nausea, and vomiting. They also said that the outbreak does not mean that the ship is unsafe and that there are no reports of anyone with fevers or any other […continues] 110010 101001 111100 001101 110010 101001 111100 001101
… Retail is another major employer in the county, with 5,280 people employed in April 2010. Nearly \$2 billion annually is spent on retail purchases in the city. The Uptown Meridian offers over one hundred shopping venues, including department stores, specialty shops, restaurants, eateries, and United Artists Theatres. Phase I of the construction of Meridian Crossroads, a shopping center in the Bonita Lakes area, was completed in November 2007, providing a major boost to retail in the area. Also, the shopping district on North Hills Street has continued to expand, and in March 2007, additional retail and office space was opened […continues] Unprecedented investments have been made recently to revitalize downtown Meridian, with more than a hundred new condominiums and commercial developments and \$250 million invested since 2013 to build retail spaces and renovate properties. As of 2019, the total sales tax revenue for the county […continues] 00001000 11001000 11111011 01000010 00001000 11001000 11111011 01000010
… The episode was directed by Andrew Overtoom for the animation, and was written by Walt Dohrn, Paul Tibbitt, and Merriwether Williams. Dohrn and Tibbitt also served as storyboard directors, and Carson Kugler, William Reiss, and Erik Wiese worked as storyboard artists. It originally aired on Nickelodeon in the United States on September 21, 2001. In this episode, SpongeBob reads a "bad word" off a dumpster behind the Krusty Krab, but does not know what it means. Patrick explains to him that it is a "sentence enhancer" which is used […continues] The main plot revolves around Patrick’s decision to give up his nickname, "Patrick Star", as he believes it does not accurately reflect his personality. As a result of this decision, he becomes increasingly obsessed with gaining a new nickname, leading several of the show’s characters […continues] 00001000 11001000 11111011 01000010 00001000 11001000 11111011 01000010
@@ Appendix B Some Experimental Results
Appendix C Pseudo-code of our procedures
Algorithm 3 illustrates our procedure of selecting the optimal Reed-Solomon code scheme. Algorithm 4 illustrates our procedure for detecting whether a piece of text is watermarked by our watermark algorithm. Algorithm 5 illustrates our algorithm for computing the provable bound for each watermarked paragraph. Algorithm 6 illustrates our algorithm for computing the $\mathsf{PI}$ array used in Algorithm 5.
@@ Appendix C Pseudo-code of our procedures
Algorithm 4 Detect zero-bit watermark
  Input: Text length $T$; text sequence $S_{0}\cdots S_{T-1}$; vocabulary set $V$; Reed-Solomon code scheme $(\hat{n},\hat{k},\hat{t})_{2^{\hat{m}}}$; pseudo-random token-to-segment mapping $M:[0,|V|-1]\rightarrow[0,\hat{n}-1]$; hash function $\mathsf{hash}$; secret key $sk$; threshold Z
  Output: True or False
  Initialize $\mathsf{COUNT}$ as $\hat{n}\times 2^{\hat{m}}$ matrix with 0
  for $i=0,1,\cdots,T-1$ do
     Compute segment index $p=M[S_{i-1}]$
     for $j=0,1,\cdots,2^{\hat{m}-1}-1$ do
        Compute random seed $s=\mathsf{hash}(sk,S_{i-1},j)$
        Use seed $s$ to select set $Q\subset V$, where $|Q|=\frac{|V|}{2}$
        if $S_{i}\in Q$ then
           $\mathsf{COUNT}[p][2j]\mathrel{+}=1$
        else
           $\mathsf{COUNT}[p][2j+1]\mathrel{+}=1$
        end if
     end for
  end for
  Initialize $sum=0$
  for $p=0,1,\cdots,\hat{n}-1$  do
     $sum\mathrel{+}=\max_{j\in[0,2^{\hat{m}}-1]}\mathsf{COUNT}[p][j]$
  end for
  Calculate z value: $z=\frac{sum-\frac{1}{2}T}{\frac{1}{2}\sqrt{T}}$
  return $z>Z$
@@
Appendix D Real-World User Generated Text Tracing
Table 6: Real-world user generated text tracking
Precision Recall User match rate
Text from Service (%) 99.6 100 97.6
We also evaluate our watermarking method under a setting that is close to real-world AI-generated text auditing applications: we consider a chatbot service with 1 million different users. There are many pieces of text of interest distributed over the Internet, with a proportion of them generated by this service and others by human written or by other LLM services. The service provider wants to be able to distinguish which pieces of text are generated by their service, and concurrently traces the original user associated with each of these texts.
For this experiment, we assume there are 750 texts, where 250 are human-written natural text, 250 are generated by OpenAI, and 250 are generated by their service. All of the texts have a length of approximately 200 tokens.
As $5762^{20}=1,048,$ 576, we can use a 20-bit string to represent a user’s unique ID. Thus, we use our Algorithm 1 to embed 20 bit long ID of each user into the text generated by this service. For other hyperparameters, we all follow the default experiment settings. For detection of service generated text we adopt Algorithm 4, and use threshold $Z=8$.
The experiment’s results are shown in Table 6. Our detection method achieves high precision and recall on discerning service generated text and other text. Note that here, other text includes both human-written text and OpenAI generated text. This shows that our detection algorithm not only can distinguish service generated text from human-written text, but also can distinguish service generated text from text generated by other LLMs.
After our detection method accurately identifies service generated text, our watermark’s extract functionality has strong performance. Achieving a user match rate of nearly $98\%$ is remarkably high considering there are 1 million users as possible candidates. Randomly guessing can only achieve $0.0001\%$ match rate under this setting.
Appendix E Analysis of Boroujeny et al. [6]
@@ Appendix D Real-World User Generated Text Tracing
Appendix F Differences between Match Rate and Bit Accuracy
In our work we focus on using match rate as our evaluation metric for watermark correctness. Several previous work [40, 66, 67, 68] in watermarking adopted another metric, bit accuracy (the ratio of correctly extracted bits). However, in many cases, the gap between match rate and bit accuracy may be large. For example, if we use Yoo et al. [68] to embed 20 bits into a 200 length paragraph, it achieves an bit accuracy of 94.7%, while the match rate is only 49%. Consequently, achieving high bit accuracy does not necessarily ensure a correspondingly high match rate. In practical applications, where the extracted message is critical, even minor extraction errors can be unacceptable. Thus, using the match rate as a metric provides a better reflection of performance in real-world scenarios.

### v5

@@ 5.1 Experiment setup
Metrics and evaluation methods. We mainly use match rate to measure the proportion of generated text that can exactly extract embedded watermark message without error. We also test bit accuracy (the ratio of correctly extracted message bits) which is adopted in several previous work [42, 69, 70, 71]. We use edit distance to quantify the attacker’s change to the watermarked text. Edit distance (Levenshtein distance) [47] measures the difference between two texts defined as the minimum number of single-token edits (insertions, deletions, or substitutions) required to change one text into the other. Additionally, to measure theoretical robustness, we define the metric robust bound for a dataset as the average robust bound for each generated text. The robust bound for each paragraph is calculated using Algorithm 4 in Appendix B. We set the error probability threshold in Definition 4.2 as $\alpha=10^{-3}$. For text quality, following the work in [33], we use perplexity (PPL) computed by a large oracle model GPT-3.5 (text-davinci-003) [52].
5.2 Main results
Table 2: Compare our method with existing baselines, on match rate and extraction time (per text paragraph). Match rate and bit accuracy are shown in percentages. "NA" means running the corresponding experiment on the dataset would take a few weeks or months, so we decided not to proceed with it.
@@ 5.2 Main results
Match
rate
Bit
Acc
Time
Match
rate
Bit
Acc
Time
Match
rate
Bit
Acc
Time
Match
rate
Bit
Acc
Time
Match
rate
Bit
Acc
Time
Fernandez et al. [19] 99.6 100.0 0.12 99.6 100.0 0.34 99.2 99.9 5.04 98.0 99.8 110 NA NA
29000
Wang et al. [65] 99.6 100.0 0.16 98.8 99.9 0.58 98.4 99.8 3.14 97.2 99.6 35.5 NA NA
8300
Yoo et al. [71] 86.4 96.5 0.01 73.6 94.6 0.01 49.2 90.8 0.01 30.4 89.4 0.01 8.4 81.2 0.01
Cohen et al. [13] 93.2 98.7 0.01 88.8 97.0 0.02 78.4 95.3 0.02 65.6 94.7 0.03 27.2 89.7 0.04
Ours 98.8 99.9 0.04 98.0 99.7 0.06 97.6 99.6 0.10 96.0 99.5 0.18 94.0 99.1 0.6
@@
Table 3: The impact of different Reed-Solomon code parameters. The scheme adopted by our watermark is in red.
Bit length 16 20 24 32
RS Code $(6,4,1)_{2^{4}}$ $(8,4,2)_{2^{4}}$ $(4,2,1)_{2^{8}}$ $(6,4,1)_{2^{5}}$ $(8,4,2)_{2^{5}}$ $(7,5,1)_{2^{4}}$ $(5,3,1)_{2^{8}}$ $(7,3,2)_{2^{8}}$ $(6,4,1)_{2^{6}}$ $(6,4,1)_{2^{8}}$ $(8,4,2)_{2^{8}}$ $(10,8,1)_{2^{4}}$
Match rate 98.0 96.4 97.2 97.6 95.6 97.6 96.0 94.4 95.6 94.0 91.2 90.4
Table 4: User study on evaluating watermarked text quality.
Avg Score STD
Unwatermarked 3.816 0.543
Watermarked 3.844 0.562
@@
Our method maintains the quality of texts generated by LLMs. We also explore the influence of our watermark on the quality of text generated by an LLM. Experiments are conducted on OpenGen dataset using 3 different models—Llama2, Guanaco, Falcon, and different bit lengths of 12, 16, 20, 24, 28, and 32. Other hyperparameters are under the default settings. Figure 5 uses box plots to compare the perplexity distribution of human-written text with that of texts generated by three different models (Llama2, Guanaco, and Falcon), both with and without watermarking. Note that a lower perplexity indicates higher text quality. The results indicate that watermarking only slightly harms the quality of AI-generated text, but the effect is minor. Prior to the application of our watermark, the perplexity values associated with text generated by Falcon predominantly fall within the interval of $[6,10]$. Following the integration of our watermark, a very slight elevation in perplexity occurs, with the generated text now primarily falling within the range of $[6.5,10.5]$. The results suggest that our watermarking method subtly affects the quality of the generated text, ensuring the utility of watermarked AI-generated content in practical applications.
We also conduct a user study to assess the text quality following [74]. We generate 100 watermarked text and 100 unwatermarked text using LLaMA2-7B model on the OpenGen dataset, and ask 25 participants to rate the text quality on a 5-points Likert Scale from 1 (poor) to 5 (excellent). Each generated text receives 2 rated scores. The average score and standard deviation are presented in Table 4. The results indicate that, the text quality degradation caused by the watermark is also subtle according to human evaluation.
Our method outperforms baselines [71, 19, 65]. To compare our approach with baselines [71, 19, 65, 13], we measure match rate and extraction time with bit lengths set to 12, 16, 20, 24, 32 separately, while other parameters are kept at their default values. We do not compare with  [7] since they are not open-sourced. Additionally, their method is not practical when embedding 20-bit messages, see our analysis in Appendix C.
@@
Impact of Reed-Solomon code parameters. To demonstrate the effectiveness of our ECC scheme selection algorithm, we evaluate different RS code parameters under different bit lengths. The results are presented in Table 3. Across the bit lengths of 16, 20, 24, 32, our selected RS code consistently achieves the highest match rate for our watermark. It can also be observed that increasing the number of recovery bits often fails to improve the match rate. This is because although more recovery bits increase the recover ability of ECC, it also results in larger $n$ which exacerbates the imbalance in token assignment. Although our algorithm generally selects the ECC with the smallest $n$, it does not consistently do so. For example, for bit length $16$, RS code $(6,4,1)_{2^{4}}$ chosen by our algorithm is better than $(4,2,1)_{2^{8}}$. This is because compared to embedding $6\times 4=24$ bits with $(6,4,1)_{2^{4}}$, $(4,2,1)_{2^{8}}$ requires embedding $4\times 8=32$ bits into generated text, resulting in insufficient text entropy to encode all 32 bits, thereby increasing the likelihood of extraction errors. In conclusion, our ECC code selection algorithm simultaneously considers the number of segments $n$, the code rate, and recovery rate to achieve the best performance.
@@ 6 Conclusion
7 Acknowledgment
We would like to extend our gratitude to the reviewers and shepherd for their constructive feedback on our work. We are also grateful to Polyhedra for providing the computing resources necessary for our experiments. Additionally, we thank Jian Liu, Zixin He for their valuable feedback.
@@ 7 Acknowledgment
Appendix B Pseudo-code of our procedures
Algorithm 3 illustrates our procedure of selecting the optimal Reed-Solomon code scheme. Algorithm 4 illustrates our algorithm for computing the provable bound for each watermarked paragraph. Algorithm 5 illustrates our algorithm for computing the $\mathsf{PI}$ array used in Algorithm 4.
@@ Appendix B Pseudo-code of our procedures
Appendix C Analysis of Boroujeny et al. [7]